OpenStudy (anonymous):

NEED HELP !!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2) ? A. y = 1/3x - 6 B. y = 1/3x -14 C. y = 3x -6 D. y = 3x - 14

OpenStudy (anonymous):

turn 3x - y = -3 to y = Mx+b m is the slope use y-y1 = m(x+x1) plug (4, -2) the turn it in to slope intercept form of y =mx+b Hint **** parallel line has the same slope

OpenStudy (anonymous):

Thanks